# Spectral-Spatial Transformer with Active Transfer Learning for Hyperspectral Image Classification

### v1

@@ V Active Transfer Learning Effects on the same dataset
Strawberry 95.1180 95.7662 97.8227 98.6230 99.0164 99.5082
Cowpea 84.8290 93.2934 96.1501 97.7410 98.7254 99.0419
Soybean 74.2951 87.1475 93.3696 96.9861 98.2694 99.2416
Sorghum 96.9742 96.8248 98.0201 98.9914 99.5890 99.7011
Water spinach 30.3333 86.5 89.3333 93.3333 98.8333 100
Watermelon 46.3812 70.9620 79.4792 87.4669 93.2480 95.5428
Greens 71.4769 85.7384 91.6666 94.5121 95.9688 96.8157
Trees 72.1659 85.3153 90.9222 94.5933 96.0952 97.3968
Grass 70.2788 91.5504 95.4795 97.4651 98.4157 98.9860
Red roof 89.4256 95.3974 98.4214 99.4104 99.6006 99.7907
Gray roof 92.8799 94.8078 96.6410 98.2022 98.9355 99.4795
Plastic 24.5108 63.5869 80.5978 92.8804 95.5978 97.5543
Bare soil 60.6186 72.1588 81.8560 87.1434 91.5313 94.0105
Road 84.4504 86.8426 93.2758 97.0689 98.9224 99.3642
Bright object 27.1126 69.8943 92.1849 83.2746 88.7323 95.4225
Water 98.9098 98.6312 99.4694 99.6870 99.8435 99.8992
@@
[…]
Figure 8: HC Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table 8.
@@
Figure 8 illustrates the OA, AA, and $\kappa$ as functions of training sample size across six iterations. The trends clearly demonstrate that increasing the number of training samples significantly improves all three metrics, with the model achieving near-optimal performance as the sample count reaches 13,532. The steep initial rise in accuracy, particularly for AA, indicates that even a moderate increase in sample size enables the model to generalize better across all classes. $\kappa$, which measures agreement beyond chance, also exhibits steady growth, confirming improved classification consistency. Figure 9 showcases the ground truth maps for each training sample size, visually representing the model’s learning progress. With 1,287 samples, the classification appears fragmented, reflecting limited learning. As the sample size increases, the maps exhibit progressively finer details, and class boundaries become more accurate and well-defined. By the final iteration with 13,532 samples, the map closely approximates the true spatial distribution, demonstrating the model’s capability to effectively leverage larger datasets for classification.
TABLE VI: HH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration.
Class Per-Class Accuracy for Training Samples
@@
Red roof 94.8433 97.7635 98.0056 99.5868 99.7435 99.7578
Road 70.6719 85.5353 91.7425 94.4191 96.0706 95.1025
Bare soil 92.6496 95.4999 97.1038 97.8828 99.0651 99.0926
Cotton 99.3140 99.2087 99.4794 99.4181 99.8873 99.7133
Cotton firewood 46.1241 62.3673 79.0286 92.2161 95.1109 98.4239
Rape 93.5050 95.1658 97.9397 99.0349 99.4120 99.3222
Chinese cabbage 81.3806 93.4782 96.5482 96.5648 98.8217 98.3820
Pakchoi 36.2604 67.8835 81.3024 89.1958 93.7839 96.4479
Cabbage 93.8990 97.1528 99.1865 99.7041 99.6487 99.4638
Tuber mustard 68.4524 83.8631 92.4640 94.9491 97.3535 96.1271
Brassica parachinensis 71.7087 90.5574 94.4252 94.8792 98.4746 97.9299
Brassica chinensis 44.4047 72.9953 87.9606 96.0464 96.3368 98.2800
Small Brassica chinensis 69.7174 86.8757 93.7799 94.5352 97.8141 97.7963
Lactuca sativa 73.4638 86.4328 95.8945 97.1995 98.6677 98.6677
Celtuce 59.4810 91.6167 96.4071 97.2055 96.8063 97.0059
Film covered lettuce 88.2401 95.2630 98.4301 98.1823 99.6144 97.6590
Romaine lettuce 82.5249 86.8438 94.9501 99.4019 98.6046 99.8671
Carrot 74.3781 94.4651 95.0248 95.2114 98.0721 99.0049
White radish 85.2617 91.8273 96.5564 95.9825 98.4848 98.7144
Garlic sprout 58.5771 88.6976 91.9678 95.0659 97.8772 94.1480
Broad bean 13.8554 53.7650 71.2349 79.0662 94.2771 94.8795
Tree 66.2871 90.9900 92.3267 97.9702 99.1584 98.8118
@@
[…]
Figure 10: HH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table VI.
@@
Table VI illustrates the per-class accuracy for a variety of land cover categories as the number of training samples rises from 1933 to 20,322. Significant accuracy gains are observed in challenging classes such as Cotton firewood, Pakchoi, and Broad bean, where accuracy improved from 46.12% to 98.42%, 36.26% to 96.44%, and 13.85% to 94.88%, respectively. These improvements highlight the model’s ability to learn complex patterns with more training data. In contrast, classes with inherently distinct features, such as Cotton and Red roofs, already achieved high accuracy with fewer samples and exhibited marginal improvements. Table VI also includes computational metrics such as training and testing times and the total number of FLOPs and parameters. Despite increasing data, the training and testing times remain relatively stable, showcasing the model’s computational efficiency. For example, training times range from 272.05 seconds to 291.68 seconds, and testing times fluctuate minimally between 22.42 seconds and 23.42 seconds. On the aggregate level, the OA, AA, and $\kappa$ demonstrated consistent growth across iterations, as shown in Figure 10. For instance, OA increased from 87.83% to 98.97%, while AA rose from 71.14% to 97.94%, reflecting the model’s robust performance across diverse classes. The $\kappa$ followed a similar upward trajectory, indicating reduced misclassification rates and alignment with ground truth. Figure 11 presents ground truth maps corresponding to the training sample counts. These visualizations show the gradual enrichment of spatial details and class boundaries as training data increases. The progression from 1933 to 20,322 samples demonstrates the model’s growing ability to capture intricate spatial and spectral variations. Notably, the maps for higher sample counts, such as 16,791 and 20,322, closely resemble the actual ground truth, highlighting the efficacy of the dataset’s size in improving model generalization.
TABLE VII: QUH Dataset: Per-class accuracy for training samples, showing accuracy improvements as the sample count increases per iteration.
Class Per-Class Accuracy for Training Samples
@@
Trees 91.8324 97.0037 97.3201 98.8538 99.0666 99.3291
Concrete building 92.0907 97.9266 98.9772 98.8401 99.5365 99.6869
Car 18.3401 81.5728 90.1334 92.7742 91.4828 96.6773
Ironhide building 98.7302 98.9350 96.5185 99.4675 99.5699 99.6108
Plastic playground 93.9587 97.1718 98.5064 98.9886 99.1788 99.6509
Asphalt road 91.9654 94.5551 96.3140 98.5950 98.6989 98.7598
@@
[…]
Figure 12: QUH Dataset: OA, AA, and $\kappa$ accuracy as a function of training samples in each iteration as shown in Table VII.
@@
Table VII presents the per-class accuracy of the QUH dataset across six iterations of training samples, highlighting the progressive improvement in accuracy as the number of samples increases. Classes such as ”Trees” and ”Concrete Building” exhibit significant gains, with accuracies exceeding 99% in the final iterations. Notably, the ”Car” class starts with lower accuracy (18.34%) due to limited initial sample representation but shows remarkable improvement, reaching 96.68% by the last iteration. The table also reports training and testing times, which grow modestly with larger sample counts, and provides the model’s computational metrics, such as FLOPs and parameters, indicating efficient scaling. Figure 12 visualizes the OA, AA, and $\kappa$ as functions of training sample size, consistent with the trends in Table VII. All three metrics show steady improvement, converging towards optimal values as sample size increases, underlining the model’s robustness and scalability. Figure 13 illustrates ground truth maps for the dataset, corresponding to each training sample size. These maps visually demonstrate the enhancement in spatial coverage and label consistency as the number of training samples increases, corroborating the quantitative improvements observed in Table VII and Figure 12.
VI Active Transfer Learning Effects on Across dataset
@@
This section explores the impact of active transfer learning (ATL) in enabling effective across-dataset classification. By leveraging the actively trained spatial-spectral transformer (SST) on one dataset and fine-tuning it with a limited fraction of samples from other datasets, we assess its ability to generalize and adapt to varying spectral-spatial distributions. The results highlight the significance of ATL in achieving high accuracy with minimal fine-tuning across diverse datasets.
Table VIII presents the OA, AA, and $\kappa$ coefficient for cross-dataset classification, where the ATL-SST is initially trained on the SA dataset and subsequently fine-tuned using 10% of the samples from other datasets (PU, UH, HC, HH, and QUH). The results indicate that the model achieves high accuracy across datasets, with $\kappa$ values consistently above 93%, demonstrating effective transfer learning. Training and testing times vary significantly across datasets due to differences in data volume and complexity, with QUH requiring the longest computation times. Figure 14 provides visual representations of classification outcomes, showcasing the trained SST on SA and its fine-tuned performance on PU, HC, HH, and UH datasets. Each subfigure illustrates distinct patterns, emphasizing the model’s adaptability to different data distributions. Together, the table and figure underscore the SST’s robust cross-dataset generalization capabilities, particularly with minimal fine-tuning.
@@
Table IX and Figure 15 illustrate the performance of the SST trained on the PU dataset and subsequently fine-tuned and evaluated on other datasets, including SA, UH, HC, and HH. Table IX provides detailed metrics—highlighting the transferability and generalization ability of the SST model. The actively trained SST on PU achieves a $\kappa$ of 98.51% and an OA of 98.88% on its test data, demonstrating robust performance on the source dataset. When fine-tuned with only 10% samples from other datasets, the model maintains competitive accuracy, with OA values exceeding 96% for most target datasets. Training and testing times are also reported, indicating computational efficiency during model fine-tuning and evaluation. While the training time scales with dataset complexity, testing times remain relatively low, demonstrating the efficiency of the SST architecture. Figure 15 visually complements these results by showcasing the SST model’s classification maps on the source dataset (PU) and various target datasets (SA, HC, HH, and UH). The maps indicate a strong generalization capability, as the fine-tuned model effectively captures spectral-spatial features across diverse datasets with minimal sample utilization. This analysis underscores the SST’s potential for cross-dataset classification, even with limited target domain data.
@@
Table X and Figure 16 present the performance of the spatial-spectral transformer (SST) actively trained on the UH dataset and fine-tuned with 10% samples from other target datasets, including SA, PU, HC, and HH. Table X outlines key metrics—highlighting the SST model’s capability to generalize and adapt to new datasets with minimal fine-tuning. The actively trained SST on UH achieves high baseline performance, with a $\kappa$ of 97.38% and an OA of 97.58% on its test data. When fine-tuned with samples from other datasets, the model demonstrates strong transferability, achieving an OA of 98.47% on SA and 97.75% on PU. Although the performance is slightly reduced for HC and HH, the OA values remain above 94%, underscoring the model’s robustness. Training and testing times scale with dataset complexity, with SA and PU requiring less computation compared to HC and HH, as seen in the recorded durations. Figure 16 complements the table by visualizing classification results. The maps indicate that the fine-tuned SST effectively captures spatial and spectral features across diverse datasets. Specifically, the SST retains consistent classification quality on target datasets despite using only 10% of the samples for fine-tuning. This outcome demonstrates the SST’s adaptability and its potential for cross-dataset applications in classification.
@@ VII Comparison with State-of-the-art Methods
The following Tables and Figure illustrate a comparative evaluation of various HSI classification methods on the University of Houston (UH), Pavia University (PU), and Salina (SA) datasets. The methods compared include Attention Graph CNN (AGCN), Hybrid Spatial-Spectral Transformer (HybViT), Spatial-Spectral Mamba (SSMamba), Spatial-Spectral Transformer (SST), Wavelet-Based Spatial-Spectral Transformer (WaveFormer), Pyramid-Based Spatial-Spectral Transformer (PyFormer), and Wavelet-Based Spatial-Spectral Mamba (WaveMamba) [50]. These state-of-the-art approaches are benchmarked against the proposed method, ATL-SST, under consistent training, validation, and testing configurations, as summarized in the Tables.
TABLE XI: UH dataset: Comparison of various HSI classification methods (Attention Graph CNN = AGCN, Hybrid Spatial-spectral Transformer = HybViT, Spatial-spectral Mamba = SSMamba) highlighting performance metrics.
@@
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
43 582 626 95.3674 93.6102 94.5686 94.8881 85.3035 91.8530 93.9297 98.4025
42 585 627 99.8405 99.3620 99.2025 99.0430 100 96.6507 98.4051 99.5215
26 323 348 99.1379 100 99.7126 98.5632 95.6896 93.6781 97.7011 100
44 578 622 94.6945 97.5884 97.2668 97.5884 98.8745 96.6237 94.8553 97.9099
42 579 621 100 99.8389 100 100 94.5249 96.6183 96.9404 100
17 146 162 87.0370 80.2469 98.1481 88.8888 88.8888 80.8641 87.6543 98.1481
106 528 634 89.7476 91.9558 92.2712 94.4794 89.2744 80.4416 85.4889 94.4794
81 541 622 87.2990 90.3536 92.7652 96.4630 87.6205 83.9228 90.9967 95.9807
81 545 626 69.9680 91.2140 93.7699 90.0958 71.7252 65.0159 83.2268 97.9233
69 544 614 96.0912 96.7426 99.5114 99.0228 84.8534 80.2931 89.7394 100
70 547 618 94.6601 95.9546 96.7637 94.8220 90.9385 86.0841 92.8802 99.0291
68 548 617 81.0372 99.0275 98.8654 97.4068 80.8752 92.7066 92.5445 97.5688
55 180 234 87.6068 82.0512 79.4871 72.2222 57.6923 34.6153 73.5042 77.7777
15 199 214 95.3271 96.2616 97.1962 97.1962 97.1962 92.5233 98.5981 100
28 302 330 94.5454 95.4545 100 100 100 81.5151 95.7575 97.8787
@@
$\kappa$ 90.6479 94.6165 96.0141 95.3937 87.6230 84.2658 91.1925 97.3807
OA 91.3506 95.0232 96.3140 95.7418 88.5562 85.4690 91.8562 97.5781
AA 91.4906 93.9775 95.9686 94.7120 88.2305 83.5604 91.4815 96.9746
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba
@@
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
422 2893 3316 99.4571 97.1351 97.4065 97.5572 96.8636 88.7515 97.6477 98.7937
535 8790 9324 98.0158 99.1419 99.0240 99.5388 98.9382 94.7447 99.5388 99.9034
218 832 1049 69.0181 84.6520 84.7473 82.9361 93.5176 74.7378 75.9771 91.3250
161 1371 1532 97.5195 96.2793 96.0182 95.1697 97.3237 92.2323 94.3864 99.5430
15 657 673 97.7711 100 100 100 100 98.9598 98.2169 100
311 2204 2514 99.2044 97.7326 97.1360 96.2609 94.3516 93.3571 98.6077 99.6420
114 551 665 87.9699 88.1203 92.9323 91.2781 75.0375 82.2556 92.3308 99.5488
405 1436 1841 93.9163 89.5708 90.0597 92.2868 63.0635 81.9663 88.4845 95.8718
64 409 474 97.0464 93.4599 96.8354 94.9367 89.4514 89.4514 90.5063 98.9451
@@
$\kappa$ 95.0080 95.3372 95.5679 95.6819 91.6276 88.1644 94.9488 98.5130
OA 96.2268 96.4840 96.6570 96.7505 93.6880 91.0183 96.1941 98.8778
AA 93.3243 94.0102 94.9066 94.4405 89.8386 88.4952 92.8551 98.1747
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba
@@
As shown in Table XI, the results highlight the superior performance of ATL-SST across most experimental settings. For OA, ATL-SST consistently achieves the highest values, with a maximum of 100% in multiple scenarios, underscoring its robustness and effectiveness. Similarly, ATL-SST outperforms all comparative methods in terms of $\kappa$ and AA, achieving top scores of 97.38% and 96.97%, respectively. This reflects the method’s ability to effectively classify diverse classes within the hyperspectral dataset. Notably, ATL-SST demonstrates significant improvement over earlier methods like AGCN and WaveMamba, with AGCN yielding lower performance metrics (e.g., 91.35% OA and 90.64% $\kappa$) in many cases. While HybViT and SST exhibit strong results, particularly in individual settings, their performance is generally less consistent compared to ATL-SST. The training time analysis also reveals ATL-SST’s computational efficiency, with an average training time comparable to SST and WaveFormer, and significantly lower than methods like WaveMamba and PyFormer. This balance of performance and efficiency makes ATL-SST a promising solution for practical HSI classification applications. The ground truth maps preseneted in Figure 17 visually depicts the classification results for the UH dataset. It is evident from the maps that ATL-SST provides more precise and coherent classification results, with fewer misclassified pixels and better-defined class boundaries compared to the comparative methods. For instance, methods like AGCN and WaveMamba exhibit noticeable misclassifications and noise, whereas ATL-SST maintains high spatial and spectral accuracy. The enhanced performance can be attributed to ATL-SST’s hybrid query mechanism and spatial-spectral feature integration, which enable more effective discrimination of classes within the hyperspectral data.
The results presented in Table XII and Figure 18 comprehensively compare the performance of the proposed ATL-SST method with state-of-the-art techniques for HSI classification on the PU dataset. Quantitative results in the Table highlight the superiority of ATL-SST across key performance metrics, while the qualitative analysis in the figure illustrates its effectiveness in generating accurate classification maps.
In terms of accuracy, ATL-SST consistently outperforms all other methods, achieving the highest OA of 98.8778%, AA of 98.1747%, and $\kappa$ of 98.5130%. These results are a significant improvement over the next best-performing methods, such as WaveMamba, which achieves an OA of 96.1941%. The robust performance of ATL-SST is evident across varying data splits, including scenarios with limited training samples. For instance, with only 15 training samples, ATL-SST achieves 100% accuracy, matching the best results from methods like WaveFormer while surpassing models such as AGCN, which only reaches 97.7711%. This demonstrates ATL-SST’s ability to generalize effectively across diverse data distributions and sample sizes. From a computational perspective, ATL-SST maintains competitive efficiency, with a training time of 124.78 seconds, comparable to HybViT and SST while significantly outperforming computationally intensive models like WaveMamba and WaveFormer, which require 1382.70 and 112.68 seconds, respectively.
Qualitative results presented in Figure 18 further validate the quantitative findings. The classification maps generated by ATL-SST are visually superior, exhibiting precise boundary delineation and accurate region classifications that closely match the ground truth. In contrast, competing methods, such as AGCN and SSMamba, show noticeable errors, particularly in regions with complex class boundaries or high spectral variability. Transformer-based models like HybViT and PyFormer produce smoother classifications but still fail to capture the finer details that ATL-SST excels at. While wavelet-based methods such as WaveFormer and WaveMamba improve spectral representation, their outputs show slight boundary smoothing, leading to less accurate segmentation compared to ATL-SST.
@@
Tr Va Te AGCN HybViT SST WaveFormer PyFormer SSMamba WaveMamba
41 964 1004 99.5019 99.9003 99.0039 98.5059 100 99.4023 100 100
@@
80 908 988 100 99.3927 99.5951 99.4939 100 97.3684 98.6842 100
62 635 697 96.4131 99.8565 99.7130 100 100 98.4218 99.1391 100
118 1221 1339 86.0343 99.8506 99.7012 99.1784 94.6975 98.6557 98.2076 99.4025
77 1902 1980 100 100 100 100 100 99.8484 99.8484 100
64 1725 1790 99.8882 99.8882 100 100 100 98.8268 99.2178 100
1004 4631 5636 86.1249 93.9850 94.9609 95.2803 88.2540 87.6330 94.2512 98.8644
109 2992 3102 100 100 100 100 99.9677 99.8388 100 99.9677
94 1545 1639 98.1696 99.5118 99.1458 99.7559 93.8987 97.3764 97.6205 99.9389
58 476 534 91.9475 100 98.8764 99.8127 88.9513 96.8164 98.3146 100
56 908 963 97.9231 100 99.5846 99.8961 100 98.6500 99.8961 100
54 404 458 96.7248 100 100 100 96.2882 99.1266 100 100
40 495 535 99.6261 99.6261 100 99.0654 46.1682 94.2056 98.1308 99.8130
842 2792 3634 92.9003 91.2493 90.7264 90.0385 92.4325 76.6373 91.1667 98.2388
50 854 903 99.4462 100 99.6677 100 100 99.6677 99.5570 100
Train (s) 91.0756 147.54 139.82 144.16 1469.81 194.06 1734.50 157.85
$\kappa$ 94.3561 97.2066 97.2511 97.2427 93.9481 92.8092 96.8405 99.4281
OA 94.9233 97.4912 97.5318 97.5244 94.5575 93.5451 97.1623 99.4864
AA 96.5437 98.9538 98.8109 98.8142 93.7911 96.4047 98.3771 99.7641
AttentionGCN
SST
HybViT
PyFormer
WaveFormer
SSMamba
WaveMamba

### v2

@@ V ATL-SST Effects on the same dataset
$\kappa$ 84.71 91.39 94.78 96.89 98.12 98.83
OA 86.97 92.65 95.54 97.34 98.39 99.00
AA 70.46 84.04 90.67 94.33 96.36 97.60
@@
Table VI provides computational metrics; despite increasing data, the training and testing times remain relatively stable, showcasing the model’s computational efficiency. For example, training times range from 272.05s to 291.68s, and testing times fluctuate minimally between 22.42s and 23.42s. For instance, OA increased from 87.83% to 98.97%, while AA rose from 71.14% to 97.94%, reflecting the model’s robust performance. The $\kappa$ followed a similar upward trajectory, indicating reduced misclassification rates and alignment with ground truth. Figure 6 presents ground truth maps corresponding to the training sample counts. These visualizations show the gradual enrichment of spatial details and class boundaries as training data increases. The progression from 1933 to 20,322 samples demonstrates the model’s growing ability to capture intricate spatial and spectral variations.
TABLE VI: HH Dataset: Accuracy improvements as the sample count increases per iteration.
Class Number of Training Samples
@@
$\kappa$ 84.51 92.38 96.07 97.36 98.82 98.69
OA 87.82 93.98 96.89 97.91 99.07 98.968
AA 71.13 86.73 92.80 95.62 97.86 97.93
@@
Table VII reports training and testing times, which grow modestly with larger sample counts, and provides the model’s computational metrics, such as FLOPs and parameters, indicating efficient scaling. Figure 7 illustrates ground truth maps for the dataset, corresponding to each training sample size. These maps visually demonstrate the enhancement in spatial coverage and label consistency as the number of training samples increases, corroborating the quantitative improvements observed in Table VII.
TABLE VII: QUH Dataset: Accuracy improvements as the sample count increases per iteration.
Class Number of Training Samples
@@
$\kappa$ 88.59 95.18 96.72 98.32 98.64 99.04
OA 91.41 96.35 97.52 98.73 98.97 99.28
AA 81.15 94.52 96.29 97.91 97.92 98.95
@@
VI ATL-SST Effects on Cross Datasets
This section explores the impact of ATL in enabling effective cross-dataset classification. By leveraging the actively trained SST on one dataset and fine-tuning it with a limited fraction of samples from other datasets, we assess its ability to generalize and adapt to varying spectral-spatial distributions. The results highlight the significance of ATL in achieving high accuracy with minimal fine-tuning across diverse datasets.
@@ VI ATL-SST Effects on Cross Datasets
Table VIII presents the OA, AA, and $\kappa$ for cross-dataset classification, where the ATL-SST is initially trained on the SA dataset and subsequently fine-tuned using 10% of the samples from other datasets (PU, UH, HC, HH, and QUH). The results indicate that the model achieves high accuracy across datasets, with $\kappa$ values consistently above 93%, demonstrating effective transfer learning. Testing times vary significantly across datasets due to differences in data volume and complexity, with QUH requiring the longest computation times. Figure 8 provides visual representations of classification outcomes, showcasing the trained SST on SA and its fine-tuned performance on PU, HC, HH, and UH datasets. Each subfigure illustrates distinct patterns, emphasizing the model’s adaptability to different data distributions.
Table IX and Figure 9 illustrate the performance of the SST trained on the PU dataset and subsequently fine-tuned and evaluated on other datasets, including SA, UH, HC, and HH. Table IX provides detailed metrics, highlighting the transferability and generalization ability of the SST model. The actively trained SST on PU achieves a $\kappa$ of 98.51% and an OA of 98.88% on its test data, demonstrating robust performance on the source dataset. When fine-tuned with only 10% samples from other datasets, the model maintains competitive accuracy, with OA values exceeding 96% for most target datasets. Testing times are also reported, indicating computational efficiency during model fine-tuning and evaluation. Figure 9 visually complements these results by showcasing the SST model’s classification maps on the source dataset (PU) and various target datasets (SA, HC, HH, and UH).
@@
Table X and Figure 10 present the performance of the SST actively trained on the UH dataset and fine-tuned with 10% samples from other target datasets, including SA, PU, HC, and HH. Table X outlines key metrics, highlighting the SST model’s capability to generalize and adapt to new datasets with minimal fine-tuning. The actively trained SST on UH achieves high baseline performance, with a $\kappa$ of 97.38% and an OA of 97.58% on its test data. When fine-tuned with samples from other datasets, the model demonstrates strong transferability, achieving an OA of 98.47% on SA and 97.75% on PU. Although the performance is slightly reduced for HC and HH, the OA values remain above 94%, underscoring the model’s robustness. Testing times scale with dataset complexity, with SA and PU requiring less computation compared to HC and HH, as seen in the recorded durations. Figure 10 complements the table by visualizing classification results. The maps indicate that the fine-tuned SST effectively captures spatial and spectral features across diverse datasets. Specifically, the SST retains consistent classification quality on target datasets despite using only 10% of the samples for fine-tuning.
@@ VII Comparison with State-of-the-art Methods
The following Tables and Figures illustrate a comparative evaluation of various HSI classification methods on the UH, PU, and SA datasets. The methods compared include Attention Graph CNN (AGCN) [51], Wavelet-Based Spatial-Spectral Transformer (WaveFormer) [43], Hybrid Spatial-Spectral Transformer (HybViT) [52], Pyramid-based Spatial-Spectral Transformer (PyFormer) [12], Spatial-Spectral Transformer (SST) [53], Spatial-Spectral Mamba (SSMamba) [54], and Wavelet-Based Spatial-Spectral Mamba (WaveMamba) [55]. These state-of-the-art approaches are benchmarked against the proposed method, ATL-SST, under consistent training, validation, and testing configurations, as summarized in the Tables.
TABLE XI: UH dataset: Comparison of various HSI classification methods (Attention Graph CNN = AGCN, Hybrid Spatial-spectral Transformer = HViT, Spatial-spectral Mamba = SSM, WaveFormer = WF, PyFormer = PF, WaveMamab = WM) highlighting performance metrics.
@@
Tr Va Te AGCN [51] HViT [52] SST [53] WF [43] PF [12] SSM [54] WM [55]
43 582 626 95.36 93.61 94.56 94.88 85.30 91.85 93.92 98.40
42 585 627 99.84 99.36 99.20 99.04 100 96.65 98.40 99.52
26 323 348 99.13 100 99.71 98.56 95.68 93.67 97.70 100
44 578 622 94.69 97.58 97.26 97.58 98.87 96.62 94.85 97.90
42 579 621 100 99.83 100 100 94.52 96.61 96.94 100
17 146 162 87.03 80.24 98.14 88.88 88.88 80.86 87.65 98.14
106 528 634 89.74 91.95 92.27 94.47 89.27 80.44 85.48 94.47
81 541 622 87.29 90.35 92.76 96.46 87.62 83.92 90.99 95.98
81 545 626 69.96 91.21 93.76 90.09 71.72 65.01 83.22 97.92
69 544 614 96.09 96.74 99.51 99.02 84.85 80.29 89.73 100
70 547 618 94.66 95.95 96.76 94.82 90.93 86.08 92.88 99.02
68 548 617 81.03 99.02 98.86 97.40 80.87 92.70 92.54 97.56
55 180 234 87.60 82.05 79.48 72.22 57.69 34.61 73.50 77.77
15 199 214 95.32 96.26 97.19 97.19 97.19 92.52 98.59 100
28 302 330 94.54 95.45 100 100 100 81.51 95.75 97.87
@@
$\kappa$ 90.64 94.61 96.01 95.39 87.62 84.26 91.19 97.38
OA 91.35 95.02 96.31 95.74 88.55 85.46 91.85 97.57
AA 91.49 93.97 95.96 94.71 88.23 83.56 91.48 96.97
AGCN [51]
SST [53]
HViT [52]
PF [12]
WF [43]
SSM [54]
WM [55]
@@
Tr Va Te AGCN [51] HViT [52] SST [53] WF [43] PF [12] SSM [54] WM [55]
422 2893 3316 99.45 97.13 97.40 97.55 96.86 88.75 97.64 98.79
535 8790 9324 98.01 99.14 99.02 99.53 98.93 94.74 99.53 99.90
218 832 1049 69.01 84.65 84.74 82.93 93.51 74.73 75.97 91.32
161 1371 1532 97.51 96.27 96.01 95.16 97.32 92.23 94.38 99.54
15 657 673 97.77 100 100 100 100 98.95 98.21 100
311 2204 2514 99.20 97.73 97.13 96.26 94.35 93.35 98.60 99.64
114 551 665 87.96 88.12 92.93 91.27 75.03 82.25 92.33 99.54
405 1436 1841 93.91 89.57 90.05 92.28 63.06 81.96 88.48 95.87
64 409 474 97.04 93.45 96.83 94.93 89.45 89.45 90.50 98.94
@@
$\kappa$ 95.00 95.33 95.56 95.68 91.62 88.16 94.94 98.51
OA 96.22 96.48 96.65 96.75 93.68 91.01 96.19 98.87
AA 93.32 94.01 94.90 94.44 89.83 88.49 92.85 98.17
AGCN [51]
SST [53]
HViT [52]
PF [12]
WF [43]
SSM [54]
WM [55]
@@
Tr Va Te AGCN [51] HViT [52] SST [53] WF [43] PF [12] SSM [54] WM [55]
41 964 1004 99.50 99.90 99.00 98.50 100 99.40 100 100
@@
80 908 988 100 99.39 99.59 99.49 100 97.36 98.68 100
62 635 697 96.41 99.85 99.71 100 100 98.4218 99.13 100
118 1221 1339 86.03 99.85 99.70 99.17 94.69 98.65 98.20 99.40
77 1902 1980 100 100 100 100 100 99.84 99.84 100
64 1725 1790 99.88 99.88 100 100 100 98.82 99.21 100
1004 4631 5636 86.12 93.98 94.96 95.28 88.25 87.63 94.25 98.86
109 2992 3102 100 100 100 100 99.96 99.83 100 99.96
94 1545 1639 98.16 99.51 99.14 99.75 93.89 97.37 97.62 99.93
58 476 534 91.94 100 98.87 99.81 88.95 96.81 98.31 100
56 908 963 97.92 100 99.58 99.89 100 98.65 99.89 100
54 404 458 96.72 100 100 100 96.28 99.12 100 100
40 495 535 99.62 99.62 100 99.06 46.16 94.20 98.13 99.81
842 2792 3634 92.90 91.24 90.72 90.03 92.43 76.63 91.16 98.23
50 854 903 99.44 100 99.66 100 100 99.66 99.55 100
Train (s) 91.07 147.54 139.82 144.16 1469.81 194.06 1734.50 157.85
$\kappa$ 94.35 97.20 97.25 97.24 93.94 92.80 96.84 99.42
OA 94.92 97.49 97.53 97.52 94.55 93.54 97.16 99.48
AA 96.54 98.95 98.81 98.81 93.79 96.40 98.37 99.76
AGCN [51]
SST [53]
HViT [52]
PF [12]
WF [43]
SSM [54]
WM [55]